OpenStudy (anonymous):

What is the value of the coefficient "b" when the quadratic equation y = (4x + 1)(2x - 3) is written in standard form?

OpenStudy (anonymous):

1.) Expand your equation. (4x+1)(2x-3) = 8x^2 +2x+12x -3 = 8x^2+14x-3 normal form takes the form of ax^2+bx+c, so b = 14